TWISTOR SPACES FOR RIEMANNIAN SYMMETRIC SPACES

Francis Burstall, Simone Gutt and John Rawnsley

Published as: Math. Ann. 295 (1993) 729–743

Abstract. We determine the structure of the zero-set of the Nijenhuis tensor of the natural almost complex structure J_{1} on the total space of the bundle J(G/K, g) of Hermitian structures on the tangent spaces of any even-dimensional Riemannian symmetric space G/K of compact or non-compact type.

1. Introduction

By a twistor space for a Riemannian manifold (M, g) we mean an (almost) complex manifold π: Z → M, fibred over M with complex fibres, together with some additional properties; see section 2 for the details. A basic example is the space J(M, g) consisting of all the complex structures on the tangent spaces of M which are compatible with the metric. J(M, g) has fibre the Hermitian symmetric space O(2n)/U(n) and the Riemannian connection allows this vertical complex structure on each fibre to be combined with the horizontal lift of the given complex structure on each tangent space to M to give J(M, g) a natural almost complex structure. This almost complex structure is integrable only for M conformally flat [3], and for compact symmetric spaces this means only the spheres and real projective spaces. For more general twistor spaces Z we may have integrability under weaker assumptions, so it is desirable to find such spaces.

Any twistor space with an integrable complex structure will have an image in J(M, g) which is a complex submanifold and so sits in the zero-set of the Nijenhuis tensor of the natural almost complex structure J_{1} on J(M, g). In [1], when M = G/K is an inner Riemannian symmetric space and g the invariant metric, this zero-set was shown to consist of a finite number of connected components each of which was a flag space of G fibring over G/K in a ‘minimal’ way (thus, the components were generalized flag manifolds for G compact and flag domains for G non-compact). In particular, each of these flag spaces is a twistor space. We used the property that G/K was inner (or, equivalently, that rank(G) = rank(K)) in our analysis.

It is the purpose of this note to determine the zero-set of the Nijenhuis tensor for an arbitrary even-dimensional Riemannian symmetric space. Our analysis uses similar ideas to those of [1] but takes into account the more complicated relationship between the root structure of G with respect to a maximal torus maximally embedded in K (a so-called fundamental torus) and the symmetric space structure when the space is not inner.

The main difference from the inner case results from the fact that we cannot show that Z respects the de Rham decomposition of G/K into irreducible factors. Indeed, it does

Typeset by AMS-TEX

Research partially supported by EEC contract SC1-0105-C and the FNRS of Belgium.

not, and the components of Z also turn out, in general, not to be homogeneous spaces of G. These twistor spaces appear to be new.

The components of the zero-set are expressible in terms of the τ-maximal parabolic subalgebras which were introduced in [1] where τ is the involution determining the symmetric space.

The paper is organized as follows. In section 2 we summarize the basic properties of twistor spaces for Riemannian manifolds. In section 3 we develop the properties of τmaximal parabolic subalgebras needed in the sequel. In section 4 we show that each point in the zero-set corresponds with a τ-maximal parabolic together with a certain subspace and in section 5 we show that τ-maximal parabolics determine open subsets of the zero-set which are generalized twistor spaces in the sense of [5]. In section 6 we apply our analysis of the zero-set to some examples. Example 1 looks at the Calabi-Eckmann Hermitian structures [2] on the product of two odd-dimensional spheres and shows that the images of these structures exhaust the zero-set. In example 2 we show that the Hermitian structures found by Samelson [6] also exhaust the zero-set in the case of an even-dimensional Lie group. In example 3 we apply our theory to a less familiar example and describe the zero-set for the twistor space of the symmetric space SU(2n)/Sp(n), n odd. By way of contrast with example 1, example 4 considers whether a product of odd-dimensional real Grassmannians might carry the analogue of a Calabi-Eckmann Hermitian structure. We show in theorem 6.1 that there can be no such Hermitian structures. Finally, in example 5 we apply our theory to obtain compact complex manifolds with the same fundamental group as certain compact locally symmetric spaces.

During the course of this work, we have benefited from conversations with D. Burns, J.A. Jiménez and D. Toledo. The second and third authors wish to thank the organizers of the Danish Lie Group Seminar for their hospitality at the Sandbjerg Estate. We thank the referee for some helpful remarks.

2. Generalized twistor spaces

In this section we recall some of the basic facts about twistor spaces. See [5], for more details and some examples.

Let V denote a real vector space of even dimension 2n with an inner product (·, ·). A Hermitian structure on V is an endomorphism J of V with J^{2} = −1 and which is compatible with the inner product in the sense that

(JX, JY) = (X, Y), ∀X, Y ∈ V.

We denote by J(V) the set of all Hermitian structures on V.

The orthogonal group O(V) acts transitively on J(V) by conjugation:

g · J = gJg^{−1}, g ∈ O(V).

The stabilizer at J of this action consists of elements of O(V) which are complex linear with respect to J and so is a copy of the unitary group. We denote it by U(V, J). Thus the set of all Hermitian structures on V coincides with the homogeneous space O(V)/U(V, J). This is a Hermitian symmetric space, so has an invariant complex structure which we describe next.

Denote by 𝔬(V), 𝔲(V, J) the Lie algebras of O(V) and U(V, J), respectively. The tangent space at J is isomorphic to the quotient 𝔬(V)/𝔲(V, J) which in turn can be identified with the subspace of elements of 𝔬(V) which anticommute with J. Multiplication of such elements on the left by J preserves this subspace and so induces an invariant almost complex structure on O(V)/U(V, J) which is integrable by standard results.

Let (M, g) be any 2n-dimensional Riemannian manifold. We denote by J(M, g) the bundle of all Hermitian structures on the tangent spaces of M. This is a bundle associated to the orthonormal frame bundle O(M, g) of the Riemannian metric g with fibre J(ℝ^{2n}). Since the fibre is homogeneous such an associated bundle can also be viewed as the quotient by the stabilizer: O(M, g)/U(ℝ^{2n}, J) where we pick some standard Hermitian structure J on ℝ^{2n} as a base-point. The horizontal distribution on the frame bundle coming from the Levi-Civita connection will thus descend to J(M, g) to give a horizontal distribution H. We denote by V the vertical distribution. The latter has a Hermitian structure coming from the invariant Hermitian structure on each fibre. The horizontal distribution H also has a Hermitian structure since each horizontal space H_{j} is isomorphic to T_{x}M if j is a Hermitian structure on T_{x}M. Thus j can be lifted by this isomorphism to H_{j}. We denote by J_{1} the almost complex structure on J(M, g) which we get by taking the direct sum of the natural horizontal and vertical Hermitian structures just defined.

By a twistor space for a Riemannian manifold (M, g) we mean an (almost) complex manifold π: Z → M, fibred over M with complex fibres together with some additional properties which we shall come to in a moment. If Z is a twistor space then, for x ∈ M, each z ∈ π^{−1}(x) defines a complex vector space structure j(z) on T_{x}M by identifying the latter with T_{z}Z/V_{z} where V is the vertical tangent bundle. Thus we get a map j: Z → J(M, g) (in general the j(z) are not automatically compatible with the metric g, but this is one of the extra assumptions we make).

Conversely, suppose we have a manifold Z which fibres over M with complex fibres and that we have a fibre-preserving map j: Z → J(M, g) which is holomorphic on each fibre. If we denote by V the vertical tangent bundle, as above, then the complex structure on each fibre transfers to V. We suppose we have a complement H for V, then, just as for J(M, g), each point z ∈ Z determines a complex structure on H_{z} as the horizontal lift of j(z). The direct sum of these two gives Z an almost complex structure which we also call J_{1}. If j preserves the horizontal distributions on Z and J(M, g) then it will be holomorphic with respect to J_{1} on each of these spaces by construction.

In [5] we called a manifold Z with a horizontal distribution H and such a horizontalpreserving map j: Z → J(M, g) which is holomorphic on the fibres a generalized twistor space. Clearly J(M, g) is itself a twistor space with j the identity map. As remarked in the introduction, J_{1} on J(M, g) is rarely integrable, so we look for generalized twistor spaces as possible candidates for developing Riemannian analogues of Penrose’s Minkowskian twistor theory.

3. τ-maximal parabolic subalgebras

The results in this section extend those of the appendix to chapter 4 of [1]. We assume that 𝔤 = 𝔨 + 𝔭 is the symmetric decomposition of a compact Lie algebra with respect to an involution τ and denote by suffices the intersections of subspaces of 𝔤 with 𝔨 or 𝔭.

If 𝔮 is a τ-stable parabolic subalgebra of 𝔤^{ℂ} we denote its nil-radical by 𝔫 and set 𝔩 = 𝔮 ∩ 𝔤 so that 𝔮 = 𝔩^{ℂ} + 𝔫. 𝔫 and 𝔩 are also τ-stable so we have decompositions

𝔮 = 𝔮_{k} + 𝔮_{p},  𝔫 = 𝔫_{k} + 𝔫_{p},  𝔩 = 𝔩_{k} + 𝔩_{p}.

Denote the centre of 𝔩 by 𝔷(𝔩) then we have the following definition taken from [1].

Definition 3.1. A parabolic subalgebra 𝔮 of 𝔤^{ℂ} is said to be τ-maximal if it is τ-stable and :

(i) 𝔩_{p} ⊂ 𝔷(𝔩); (ii) 𝔫 = 𝔫_{p} + [𝔫_{p}, 𝔮_{p}].

In [1] we showed how to construct τ-maximal parabolics starting from a τ-stable Borel subalgebra. Indeed, in theorem 4.29 of [1], it was shown that if 𝔟 is such a Borel subalgebra and 𝔟′ is its nilradical then 𝔟′_{p} + [𝔟′_{p}, 𝔟_{p}] is the nilradical of a τ-maximal parabolic subalgebra 𝔮. Since 𝔟′_{p} + [𝔟′_{p}, 𝔟_{p}] ⊂ 𝔟′, taking polars with respect to the Killing form gives 𝔟 ⊂ 𝔮. In fact, we also have have the converse:

Lemma 3.2. If 𝔮 is τ-maximal and 𝔟 is any τ-stable Borel subalgebra contained in 𝔮 then 𝔮 has nilradical 𝔟′_{p} + [𝔟′_{p}, 𝔟_{p}].

Proof. 𝔱 = 𝔟∩𝔤 is a τ-stable maximal toral subalgebra of 𝔤 which is contained in 𝔩 = 𝔮∩𝔤. But 𝔮 is τ-maximal so that 𝔩_{p} ⊂ 𝔷(𝔩) whence 𝔩_{p} ⊂ 𝔱 and thus 𝔩_{p} = 𝔱_{p}. Now 𝔤^{ℂ} = 𝔱^{ℂ} ⊕𝔟′ ⊕\overline{𝔟′} so that 𝔭^{ℂ} = 𝔩_{p}^{ℂ}⊕𝔟′_{p}⊕\overline{𝔟′}_{p}. On the other hand, 𝔫 ⊂ 𝔟′ so that 𝔫_{p} ⊂ 𝔟′_{p} while 𝔭^{ℂ} = 𝔩_{p}^{ℂ}⊕𝔫_{p}⊕\overline{n}_{p}. Thus 𝔫_{p} = 𝔟′_{p}. Moreover, 𝔮_{p} = 𝔩_{p}^{ℂ} ⊕ 𝔫_{p} so that we conclude that 𝔮_{p} = 𝔱_{p}^{ℂ} ⊕ 𝔟′_{p} = 𝔟_{p}. Thus [𝔫_{p}, 𝔮_{p}] = [𝔟′_{p}, 𝔟_{p}] and the result now follows immediately from the τ-maximality of 𝔮 □

Remark 3.3. In the course of the proof of lemma 3.2 we have shown that for 𝔮 τmaximal, 𝔮 ∩ 𝔭 is the 𝔭-part of a maximal toral subalgebra of 𝔤.

We now have a simple characterization of τ-maximal subalgebras given by the following theorem.

Theorem 3.4. A τ-stable parabolic subalgebra 𝔮 is τ-maximal if and only if it contains a τ-stable Borel subalgebra 𝔟 with 𝔫 = 𝔟′_{p} + [𝔟′_{p}, 𝔟_{p}].

Proof. Let 𝔮 be a parabolic subalgebra with 𝔫 = 𝔟′_{p} + [𝔟′_{p}, 𝔟_{p}] for some τ-stable Borel subalgebra 𝔟. Then 𝔟 ∩ 𝔤 is a τ-stable maximal toral subalgebra of 𝔤 which, by lemma 4.27 of [1], is fundamental. Thus theorem 4.29 of [1] says that 𝔮 is τ-maximal.

Conversely, if 𝔮 is a τ-maximal parabolic subalgebra, then it contains a τ-stable Borel subalgebra 𝔟 and Lemma 3.2 gives the required condition on its nilradical. □

4. Points in the zero-set

Let G/K be an even-dimensional Riemannian symmetric space of compact or noncompact type. The action of G as isometries on G/K lifts into J(G/K, g) and preserves Z. Since G acts transitively on G/K then Z will be G · Z_{k} = G ×_{K} Z_{k} where Z_{k} denotes the intersection of Z with the fibre of J(G/K, g) over the identity coset. If we identify the tangent space to G/K at the identity coset with 𝔭 where

𝔤 = 𝔨 + 𝔭

is the usual symmetric space decomposition of the Lie algebra 𝔤 of G, then the fibre of J(G/K, g) over the identity coset can be identified with J(𝔭), the set of all skew-symmetric transformations j of 𝔭 with j^{2} = −I. Such a transformation j has eigenvalues ±i and is determined by its +i-eigenspace which we denote by 𝔭^{+}. If 𝔤 is given an invariant bilinear form which induces the metric on G/K then 𝔭^{+} is a maximal isotropic subspace of the complexification 𝔭^{ℂ} of 𝔭. We shall use j and 𝔭^{+} interchangeably without further comment. In [1] it is shown that the condition for j to be in the zero-set of the Nijenhuis tensor is

[[𝔭^{+}, 𝔭^{+}], 𝔭^{+}] ⊂ 𝔭^{+},

or equivalently that [𝔭^{+}, 𝔭^{+}] is an isotropic subspace of 𝔨^{ℂ}.

For connected G the components of Z will have the form G · Z_{1} where each Z_{1} is a component of Z_{k}. Our goal is to describe the structure of the components of Z_{k}. Moreover, in view of the celebrated duality between symmetric spaces of compact and non-compact type, it suffices to take G compact. This is possible since, when G/K is of non-compact type, the space 𝔭^{ℂ}, the isotropic subspaces 𝔭^{+} and their K-orbits coincide with those of the compact dual U/K and thus Z_{k} is the same for both spaces.

So let 𝔤 be compact and let 𝔭^{+} be in Z_{k}. Set

𝔥 = {ξ ∈ 𝔤 : [ξ, 𝔭^{+}] ⊂ 𝔭^{+} + [𝔭^{+}, 𝔭^{+}]}

then 𝔥 is τ-stable so 𝔥 = 𝔥_{k} + 𝔥_{p} where

𝔥_{k} = 𝔥 ∩ 𝔨 = {ξ ∈ 𝔨 : [ξ, 𝔭^{+}] ⊂ 𝔭^{+}}

and

𝔥_{p} = 𝔥 ∩ 𝔭 = {ξ ∈ 𝔭 : [ξ, 𝔭^{+}] ⊂ [𝔭^{+}, 𝔭^{+}]}.

𝔥_{k} is then the Lie algebra of H_{k} = {k ∈ K : Ad_{G} k 𝔭^{+} ⊂ 𝔭^{+}}.

Lemma 4.1. 𝔥_{p} is an abelian subalgebra of 𝔤 and [𝔥_{p}, 𝔥_{k}] = 0.

Proof. Let ξ ∈ 𝔥_{k}, η ∈ 𝔥_{p} and ζ ∈ 𝔭^{+} then [η, ζ] = ∑_{i}[λ_{i}, μ_{i}] for some λ_{i} and μ_{i} in 𝔭^{+}. Thus

([ξ, η], ζ) = (ξ, [η, ζ]) = ∑_{i}(ξ, [λ_{i}, μ_{i}]) = ∑_{i}([ξ, λ_{i}], μ_{i}) = 0

so [𝔥_{p}, 𝔥_{k}] = 0. Obviously [𝔥_{p}, 𝔥_{p}] ⊂ 𝔥_{k} and if ξ, η ∈ 𝔥_{p}, ζ ∈ 𝔥_{k} then

([ξ, η], ζ) = (ξ, [η, ζ]) = 0

and hence [𝔥_{p}, 𝔥_{p}] = 0. □

Let 𝔪 denote the orthogonal complement of 𝔥_{k} in 𝔨 and 𝔪^{ℂ} its complexification.

Lemma 4.2. We have 𝔪^{ℂ} = [𝔭^{+}, 𝔭^{+}] + \overline{[𝔭^{+}, 𝔭^{+}]} and 𝔥_{k} is the centralizer of a torus in 𝔨.

Proof. Lemma 5.1 and Proposition 5.2 of [1] still apply since these are proven without the assumption that G/K is inner. □

Take a maximal toral subalgebra 𝔱_{k} of 𝔨 in 𝔥_{k}. Such a toral subalgebra exists by Lemma 4.2. Then 𝔱 = 𝔱_{k} + 𝔱_{p} is a fundamental toral subalgebra of 𝔤 where 𝔱_{p} is the centralizer of 𝔱_{k} in 𝔭. It is clear that 𝔱_{p}^{ℂ} is the zero weight space (relative to 𝔱_{k}) for 𝔭^{ℂ} as a representation of 𝔨^{ℂ} and so as of 𝔥_{k}^{ℂ}. 𝔭^{ℂ} splits into 𝔭^{+} + \overline{𝔭^{+}} as a representation of 𝔥_{k}^{ℂ} and so 𝔱_{p}^{ℂ} is the sum of the zero weight space 𝔱^{+} on 𝔭^{+} and its complex conjugate. Hence 𝔱^{+} is a maximal isotropic subspace of 𝔱_{p}^{ℂ}.

Let Δ be the root system of (𝔤^{ℂ}, 𝔱^{ℂ}) and let I denote the set of roots which vanish on 𝔱_{p}, II those which do not. If α ∈ I then the root space 𝔤_{α} lies in 𝔨^{ℂ} or 𝔭^{ℂ}. Let I_{k} and I_{p} denote the corresponding sets of roots, so Δ = I_{k} ∪ I_{p} ∪ II is a disjoint union.

Lemma 4.3. Each root of type II is non-zero on 𝔱^{+}.

Proof. The roots of a compact torus take imaginary values, so a root α of type II will be imaginary on 𝔱_{p} and hence if it vanishes on 𝔱^{+} it will vanish on the complex conjugate and so on 𝔱_{p}. This is impossible. □

For each root α choose a non-zero vector e_{α} in 𝔤_{α}. If a root α is in II then 𝔤_{α} cannot lie entirely in 𝔨^{ℂ} nor in 𝔭^{ℂ}. Thus there are non-zero elements x_{α} ∈ 𝔨^{ℂ} and y_{α} ∈ 𝔭^{ℂ} with e_{α} = x_{α} + y_{α}.

Lemma 4.4. Let 𝔭^{+} be in Z_{k}, choose 𝔱_{k}, 𝔱_{p} as above and let Δ be the roots of 𝔤^{ℂ} with respect to 𝔱^{ℂ}. Set

Φ = {α ∈ Δ : 𝔤_{α} ⊂ 𝔭^{+} + [𝔭^{+}, 𝔭^{+}]}

then Φ is closed under root addition and there exists a subset 𝔱^{+} of 𝔱_{p}^{ℂ} such that

𝔭^{+} = 𝔱^{+} + ∑_{α∈Φ∩II} ℂy_{α} + ∑_{α∈Φ∩I_{p}} 𝔤_{α}.

Proof. We examine 𝔭^{ℂ} in terms of its weight spaces as a representation of 𝔱_{k}. The zero weight space is 𝔱_{p}^{ℂ} by definition, so the zero weight space on 𝔭^{+} will be a maximal isotropic subspace 𝔱^{+} of 𝔱_{p}^{ℂ}. To finish the proof we need to show that the 𝔱_{k}-invariant complement of 𝔱^{+} in 𝔭^{+} consists of 1-dimensional weight spaces. This depends on knowing how the roots of 𝔤^{ℂ} may coincide when they are restricted to 𝔱_{k}. Obviously the restrictions of no two type I roots can coincide. Equally obviously if α is of type II then α and τα coincide if τ is the involution, but this is the only way two type II roots can coincide when they are restricted. This follows since x_{α} will be a root vector of 𝔨^{ℂ} for the restriction of α of type II. If two type II roots α, β have coincident restrictions, then both x_{α} and x_{β} would be in the same (𝔨^{ℂ}, 𝔱_{k}^{ℂ})-root space. Thus x_{α} and x_{β} are proportional, and by rescaling we can assume they are equal. Then for any ξ in 𝔱_{p} we have

α(ξ)^{2}x_{α} = [ξ, [ξ, x_{α}]] = [ξ, [ξ, x_{β}]] = β(ξ)^{2}x_{β}

so α = ±β on 𝔱_{p}. Hence α = β or α = τβ. Thus the only remaining coincidence that can happen is that the restriction of a type I root β coincides with the restrictions of a pair of type II roots α and τα.

The weight spaces for non-zero weights will be one-dimensional unless there are coincidences when roots of (𝔤^{ℂ}, 𝔱^{ℂ}) are restricted to 𝔱_{k}. By the above, weight vectors will be either type I_{p} root vectors, or the y_{α} of type II roots or a combination of these when there happens to be a coincidence for restricted roots. So suppose that β ∈ I_{p} with e_{β} ∈ 𝔤_{β} and α ∈ II and also that α = β on 𝔱_{k} with e_{β} + y_{α} ∈ 𝔭^{+}. By Lemma 4.3 we can pick an element ξ of 𝔱^{+} with α(ξ) ≠ 0 then [ξ, [ξ, e_{β} + y_{α}]] ∈ 𝔭^{+}. But this is α(ξ)^{2}y_{α} and so y_{α} ∈ 𝔭^{+}. Thus e_{β} ∈ 𝔭^{+} also.

Thus we have shown that 𝔭^{+} is composed of a maximal isotropic subspace 𝔱^{+} of 𝔱_{p}^{ℂ} together with a sum of type I_{p} root spaces and a sum of spaces of the form ℂy_{α} for type II roots α. We may now define

Φ = {α ∈ Δ : 𝔤_{α} ⊂ 𝔭^{+} + [𝔭^{+}, 𝔭^{+}]}

and we have

𝔭^{+} = 𝔱^{+} + ∑_{α∈Φ∩II} ℂy_{α} + ∑_{α∈Φ∩I_{p}} 𝔤_{α}

as required. We note that since 𝔭^{+} + [𝔭^{+}, 𝔭^{+}] is an algebra, Φ will be closed under root addition. □

With this we have the main result relating points in the zero-set of the Nijenhuis tensor to parabolic subalgebras:

Theorem 4.5. If 𝔭^{+} ∈ Z_{k} then there exists a τ-maximal parabolic subalgebra 𝔮 of 𝔤^{ℂ} with 𝔮 ∩ 𝔤 = 𝔥 and such that 𝔭^{+} = 𝔫 ∩ 𝔭^{ℂ} + 𝔥^{+} with 𝔥^{+} a maximal isotropic subspace of 𝔥_{p}^{ℂ} where 𝔥_{p} = 𝔭 ∩ 𝔮.

Proof. We use the subset Φ of the roots defined in lemma 4.4. Since 𝔭^{+} ∩ \overline{𝔭^{+}} = 0 we have Φ ∩ −Φ = ∅, whilst II ⊂ Φ ∪ −Φ. Since Φ is closed under root addition it follows that

𝔫 = ∑_{α∈Φ} 𝔤_{α}

is the nilradical of a τ-stable parabolic 𝔮 with Levi factor 𝔥^{ℂ} and that 𝔭^{+} = 𝔱^{+} + 𝔫 ∩ 𝔭^{ℂ}. Finally note that [𝔱_{p}, 𝔭^{+}] = ∑_{α∈Φ∩II} ℂx_{α} ⊂ [𝔭^{+}, 𝔭^{+}] and so 𝔱_{p} ⊂ 𝔥. Hence we must have 𝔱_{p} = 𝔥_{p}. In particular 𝔱^{+} is a maximal isotropic subspace of 𝔥_{p}^{ℂ}.

In order to see that the parabolic subalgebra 𝔮 constructed above is τ-maximal we observe that condition (i) of definition 3.1 is a consequence of lemma 4.1. Condition (ii) follows since 𝔫_{k} = [𝔭^{+}, 𝔭^{+}] by lemma 4.2. But 𝔫 is an ideal in 𝔮 so [𝔫_{p}, 𝔱_{p}^{ℂ} + 𝔫_{p}] ⊂ 𝔫 ∩ 𝔨^{ℂ} = 𝔫_{k}. □

We also have a converse to this result. Suppose we have a τ-maximal parabolic 𝔮 with Levi factor 𝔩 = 𝔮 ∩ 𝔤 then we know that 𝔩_{p} = 𝔮 ∩ 𝔭 is even dimensional and if 𝔫 is the nilradical then 𝔭^{ℂ} = 𝔫_{p} + \overline{𝔫_{p}} + 𝔩_{p}^{ℂ}. If we take a maximal isotropic subspace 𝔩^{+} of 𝔩_{p}^{ℂ} then 𝔭^{+} = 𝔫_{p} + 𝔩^{+} is maximal isotropic. In fact:

Theorem 4.6. If 𝔮 is a τ-maximal parabolic of 𝔤^{ℂ} and 𝔩^{+} is a maximal isotropic subspace of 𝔩_{p}^{ℂ} (defined as above) then 𝔭^{+} = 𝔫_{p} + 𝔩^{+} is in Z_{k}.

Proof. We have seen that 𝔭^{+} is maximal isotropic. Since 𝔫 is an ideal in 𝔮 and 𝔩_{p} is abelian then [𝔭^{+}, 𝔭^{+}] ⊂ [𝔫_{p}, 𝔫_{p} + 𝔩^{+}] ⊂ 𝔫_{k}. Further [𝔫_{k}, 𝔭^{+}] ⊂ 𝔫_{p} ⊂ 𝔭^{+}, so [[𝔭^{+}, 𝔭^{+}], 𝔭^{+}] ⊂ 𝔭^{+}. □

Consider the set Z̃_{k} consisting of pairs (𝔮, 𝔩^{+}) where 𝔮 is a τ-maximal parabolic and 𝔩^{+} is a maximal isotropic subspace of (𝔮 ∩ 𝔭)^{ℂ}. Theorem 4.5 gives us a map a: Z_{k} → Z̃_{k} a(𝔭^{+}) = (𝔥^{ℂ} + ∑_{α∈Φ} 𝔤_{α}, 𝔥^{+}) and theorem 4.6 gives us a map b: Z̃_{k} → Z_{k} defined by b(𝔮, 𝔩^{+}) = 𝔫 ∩ 𝔭^{ℂ} + 𝔩^{+} where 𝔫 is the nilradical of 𝔮.

Theorem 4.7. The maps a, b, defined above, are inverses of each other.

Proof. b∘a is clearly the identity. To see the converse, suppose we have a τ-maximal parabolic 𝔮 and a maximal isotropic subspace 𝔩^{+} of 𝔩_{p}^{ℂ} (notation as in section 3) and we set 𝔭^{+} = 𝔫_{p} + 𝔩^{+}. Take a maximal toral subalgebra 𝔱_{k} of 𝔩_{k} then 𝔱 = 𝔱_{k} + 𝔩_{p} is maximal toral in 𝔤 (see remark 3.3). Take the roots of 𝔤^{ℂ} with respect to 𝔱^{ℂ} and divide them into types I and II as usual. Type I roots vanish on 𝔩_{p}, so [𝔤_{α}, 𝔩_{p}^{ℂ}] = 0 = [𝔤_{α}, 𝔩^{+}] for α of type I. As in lemma 4.3, a root α of type II does not vanish on 𝔩^{+}, and so [𝔤_{α}, 𝔩^{+}] = 𝔤_{α} = [𝔤_{α}, 𝔩_{p}^{ℂ}]. Thus for all roots α we have [𝔤_{α}, 𝔩^{+}] = [𝔤_{α}, 𝔩_{p}^{ℂ}] and so summing over root spaces in 𝔫 we have [𝔫, 𝔩^{+}] = [𝔫, 𝔩_{p}^{ℂ}]. Intersecting with 𝔨^{ℂ} we conclude that [𝔫_{p}, 𝔩^{+}] = [𝔫_{p}, 𝔩_{p}^{ℂ}] and so [𝔭^{+}, 𝔭^{+}] = [𝔫_{p}, 𝔫_{p} + 𝔩^{+}] = [𝔫_{p}, 𝔫_{p} + 𝔩_{p}^{ℂ}] = [𝔫_{p}, 𝔮_{p}] = 𝔫_{k} since 𝔮 is τ-maximal.

This means that the 𝔥_{k} determined by 𝔭^{+} will be equal to the 𝔩_{k} of 𝔮, and so 𝔥_{p} = 𝔩_{p} and then 𝔥^{+} = 𝔭^{+}∩𝔥_{p}^{ℂ} = 𝔩^{+}. Then the 𝔭-part of the nilradical of the parabolic determined by 𝔭^{+} will be 𝔫_{p} and so we recover both 𝔮 and 𝔩^{+} from 𝔭^{+} showing that a∘b = id. □

5. The structure of the zero-set

We now associate a subset Z_{𝔮} of the zero-set of the Nijenhuis tensor of J_{1} on J(G/K, g) to the K-conjugacy class of a τ-maximal parabolic 𝔮; we continue with the notation above.

Let J(𝔩_{p}) denote the almost complex structures on the vector space 𝔩_{p} compatible with the Killing form and give J(𝔩_{p}) its natural structure of a complex manifold as in section 2. Let L_{k} be the stabilizer in K of 𝔮 in the adjoint representation of G on 𝔤^{ℂ}. Then L_{k} has Lie algebra the normalizer of 𝔮 in 𝔨. Since a parabolic subalgebra of 𝔤^{ℂ} is its own normalizer, it follows that L_{k} has Lie algebra 𝔮 ∩ 𝔨 = 𝔩_{k}. Obviously, L_{k} also preserves 𝔮_{k} and so the latter defines an invariant complex structure on K/L_{k}. Give K/L_{k} × J(𝔩_{p}) the product complex structure. Define a map ϕ: K/L_{k} × J(𝔩_{p}) → J(𝔭) by ϕ(kL_{k}, 𝔩^{+}) = Ad k(b(𝔮, 𝔩^{+})) where b is the map defined in section 4. Then we have the following proposition:

Proposition 5.1. The map ϕ defined above is holomorphic.

Proof. That the map is holomorphic follows by considering the two variables separately. The inclusion of J(𝔩_{p}) into J(𝔭) given by adding on a fixed isotropic subspace 𝔫_{p} is clearly holomorphic. Keeping the point in J(𝔩_{p}) fixed we need to see finally that the map from K/L_{k} to J(𝔭) given by conjugating a fixed element j_{0} of J(𝔭) is holomorphic. This follows from the following more general lemma.

Lemma 5.2. Let the reductive homogeneous space K/H have a complex structure given by the subspace 𝔪^{+} of 𝔪^{ℂ} where 𝔪 is the reductive summand. Let j_{0} be an H-invariant element of J(𝔭) where 𝔭 is an even-dimensional representation of K. Then the map kH ↦ k j_{0} k^{−1} is holomorphic if and only if 𝔪^{+}·𝔭^{+} ⊂ 𝔭^{+} where · denotes the infinitesimal action and 𝔭^{+} is the +i eigenspace of j_{0}.

Proof. Denote the map by ϕ and let ξ̃ denote the vector-field on K/H generated by an element ξ of the Lie algebra 𝔨 of K. Then

dϕ(ξ̃_{eH}) = [ξ·, j_{0}].

Since ϕ is equivariant it will be holomorphic if its differential at the identity coset preserves the spaces of (1,0) vectors. Thus, for ξ ∈ 𝔪^{+}, we need to have [ξ·, j_{0}] in the (1,0) space at j_{0}. The latter consists of endomorphisms A of 𝔭^{ℂ} which anticommute with j_{0} and satisfy (j_{0} − i)A = 0. Thus take ξ in 𝔪^{+} and consider

(j_{0} − i)∘[ξ·, j_{0}] = (j_{0} − i)∘(ξ·)∘j_{0} − (j_{0} − i)∘j_{0}∘(ξ·)
= (j_{0} − i)∘(ξ·)∘(j_{0} + i)

This will vanish if and only if 𝔪^{+} · 𝔭^{+} ⊂ 𝔭^{+}. □

To complete the proof of proposition 5.1, we observe that in our case 𝔪^{+} = 𝔫_{k} and 𝔭^{+} = 𝔫_{p} + 𝔩^{+} so 𝔪^{+} · 𝔭^{+} = [𝔫_{k}, 𝔫_{p} + 𝔩^{+}] ⊂ 𝔫_{p} ⊂ 𝔭^{+}. □

Theorem 4.7 shows that ϕ is injective so we may use it to view K/L_{k} × J(𝔩_{p}) as a subset of the fibre of J(G/K, g) over the identity coset and take its orbit Z_{𝔮} under G. Clearly this orbit will be in the zero-set of the Nijenhuis tensor and depends only on the K-conjugacy class of 𝔮. As a manifold it is just the homogeneous fibre bundle associated to the principal K-bundle G → G/K with fibre K/L_{k} × J(𝔩_{p}) and as such it is an example of a generalized twistor space as considered in section 2. There it is shown that such spaces have a natural almost complex structure J_{1} with respect to which the natural map to J(G/K, g) is holomorphic. In our case this map is just the inclusion map, so that we have immediately that Z_{𝔮} is an almost-complex submanifold of J(G/K, g). Since Nijenhuis tensors are natural with respect to almost-complex maps, and Z_{𝔮} is in the zero-set of the Nijenhuis tensor of J(G/K, g) it follows that the Nijenhuis tensor of J_{1} on Z_{𝔮} also vanishes and hence that J_{1} is integrable on Z_{𝔮}. We have thus shown:

Proposition 5.3. To each τ-maximal parabolic 𝔮 is associated a subset Z_{𝔮} of J(G/K, g) which lies in the zero-set of the Nijenhuis tensor of J_{1}. The latter induces an integrable complex structure on Z_{𝔮}. Z_{𝔮} depends only on the K-conjugacy class of 𝔮.

Proposition 5.4. There are only a finite number of K-conjugacy classes of τ-maximal parabolics.

Proof. Each τ-maximal parabolic 𝔮 determines a parabolic 𝔮_{k} of 𝔨^{ℂ} and there are only a finite number of K-conjugacy classes of these. It suffices to show, therefore, that the 𝔨^{ℂ} parabolic 𝔮_{k} can be contained in only a finite number of τ-maximal parabolics. But this is the case since we can choose a maximal toral subalgebra 𝔱_{k} of 𝔩_{k} which is also maximal in 𝔨. We take 𝔩_{p} to be the centralizer of 𝔩_{k} in 𝔭 (so dependent only on 𝔮_{k} and not 𝔮). Then we know 𝔱_{k} + 𝔩_{p} is a maximal toral subalgebra of 𝔤. Since a maximal toral subalgebra may only be contained in a finite number of parabolics of any kind this means that the extensions 𝔮 of 𝔮_{k} are finite in number. □

We summarize these results as

Theorem 5.5. The zero-set of the Nijenhuis tensor of J_{1} on J(G/K, g) is a finite union of complex manifolds of the form Z_{𝔮} where 𝔮 is a τ-maximal parabolic of 𝔤^{ℂ}.

6. Examples and applications

Let us see what our analysis tells us about the geometry of Z and examine some examples. First we note that the situation is rather more complicated for non-inner Riemannian symmetric spaces than for the inner spaces treated in [1]: for instance, G does not act transitively on the components of Z except when J(𝔩_{p}) is zero-dimensional, which is the case precisely when dim 𝔩_{p} = 2. We remark that dim 𝔩_{p} = rank G − rank K and so only depends on τ rather than the particular τ-maximal parabolic 𝔮. Thus G will be transitive on all components of Z if it is transitive on one.

Moreover, if G/K = G_{1}/K_{1} × G_{2}/K_{2} is an isometric splitting of G/K into a pair of evendimensional non-inner Riemannian symmetric spaces, then 𝔩_{p} = 𝔩_{p_{1}} + 𝔩_{p_{2}} but J(𝔩_{p}) ≠ J(𝔩_{p_{1}}) × J(𝔩_{p_{2}}) so that, in general, j ∈ Z_{𝔮} will not split as j = j_{1} + j_{2} with j_{i} ∈ J(G_{i}/K_{i}). Thus Z does not respect the de Rham decomposition of G/K in contrast to the case of inner symmmetric spaces (compare theorem 5.3 of [1]). Despite this, the Z_{𝔮} do not behave too badly with respect to the de Rham decomposition: if 𝔟 ⊂ 𝔤^{ℂ} is a τ-stable Borel subalgebra and

𝔤 = 𝔤_{1} ⊕ ⋯ ⊕ 𝔤_{k}

is the decomposition of 𝔤 into irreducible orthogonal symmetric Lie algebras, then it is straightforward to show that 𝔟 = 𝔟_{1} ⊕ ⋯ ⊕ 𝔟_{k} with each 𝔟_{i} a τ-stable Borel subalgebra of 𝔤_{i}^{ℂ}. Thus τ-stable parabolic subalgebras also commute with this decomposition.

Let us now consider some examples:

Example 1. Let us take our symmetric space to be a product of odd-dimensional spheres S^{2n−1} × S^{2m−1} = SO(2n) × SO(2m)/SO(2n − 1) × SO(2m − 1). In this case, rank G − rank K = 2 so that the connected components of Z are G-orbits. To find the Z_{𝔮}, we note from the above discussion that a τ-maximal parabolic subalgebra of 𝔤^{ℂ} is a sum of τ-maximal parabolic subalgebras for the factors 𝔰𝔬(2n) and 𝔰𝔬(2m) and so it suffices to find these. For this, fix x ∈ S^{2n−1} and set V = {x}^{⊥}. Let τ be the involution at x and then, under the usual identification of 𝔰𝔬(2n) with Λ^{2}ℝ^{2n}, we have as symmetric decomposition:

Λ^{2}ℝ^{2n} = Λ^{2}V ⊕ V ⊗ ℝx.

A τ-maximal parabolic is equivalent to a maximal isotropic subspace V^{+} of V^{ℂ}: we have an orthogonal direct sum

V^{ℂ} = V^{+} ⊕ V^{0} ⊕ V^{−}

with V^{±} mutually conjugate and V^{0} real and 1-dimensional and then the corresponding parabolic subalgebra 𝔮 has nilradical

(Λ^{2}V^{+} ⊕ V^{+} ⊗ V^{0}) ⊕ V^{+} ⊗ ℂx.

We note that such a parabolic determines (and is determined by) a choice of complex structure on ℝ^{2n} (equivalently, a choice of isomorphism ℝ^{2n} ≅ ℂ^{n}). Indeed, if j : ℝ^{2n} → ℝ^{2n} is a complex structure, take V^{0} = ℂjx and V^{+} to be the \sqrt{−1}-eigenspace of j on {x, jx}^{⊥}. We denote the corresponding parabolic by 𝔮_{x,j}.

Thus, if (x, y) ∈ S^{2n−1} × S^{2m−1} and τ is the involution at (x, y) then any τ-maximal parabolic is of the form 𝔮_{x,j} ⊕ 𝔮_{y,k} with j, k complex structures on ℝ^{2n} and ℝ^{2m} respectively. The 𝔭-part of the Levi-factor is then given by

𝔩_{p} = ℝjx ⊗ x + ℝky ⊗ y

giving just two choices for 𝔩^{+}: ℂ(jx ⊗ x ± \sqrt{−1}ky ⊗ y).

Observe that fixing j, k and a choice of 𝔩^{+} while letting x and y vary gives rise to a globally defined section of Z—this section is easily checked to be an integrable Hermitian structure on S^{2n−1} × S^{2m−1} and is that discovered by Calabi-Eckmann [2].

From lemma 5.4 of [1], it is known that any integrable Hermitian structure on an even dimensional manifold M, when viewed as a section of J(M), has image in Z. In the case at hand then, we conclude from the above development that for S^{2n−1} × S^{2m−1}, Z is exhausted by the images of globally defined Hermitian structures.

Example 2. Let G be an even-dimensional compact semisimple Lie group. We view G as a symmetric G × G-space G ≅ (G × G)/ΔG. The involution at the identity coset is then τ : (x, y) ↦ (y, x) so the the symmetric decomposition has

𝔨 = Δ𝔤, 𝔭 = {(ξ, −ξ): ξ ∈ 𝔤}.

Now a τ-stable Borel subalgebra of 𝔤^{ℂ} ⊕ 𝔤^{ℂ} is of the form 𝔟 ⊕ 𝔟 with Levi-factor 𝔱^{ℂ} ⊕ 𝔱^{ℂ} for a Borel subalgebra 𝔟 of 𝔤^{ℂ} and 𝔱 = 𝔟 ∩ 𝔤. We now apply theorem 3.4 to find the τ-maximal parabolic subalgebras: let 𝔟′ be the nilradical of 𝔟 and 𝔫 that of 𝔟 ⊕ 𝔟. We have

𝔫_{p} = {(ξ, −ξ): ξ ∈ 𝔟′} (𝔟 ⊕ 𝔟)_{p} = {(η, −η): η ∈ 𝔟},

so that

[𝔫_{p}, (𝔟 ⊕ 𝔟)_{p}] = Δ[𝔟′, 𝔟] = Δ𝔟′ = 𝔫_{k}.

From this we see that the τ-maximal parabolic containing 𝔟 ⊕ 𝔟 is 𝔟 ⊕ 𝔟 itself, that is, the τ-maximal parabolics are precisely the τ-stable Borels.

We now use projection onto the first factor to identify 𝔭 with 𝔤 and conclude that 𝔭^{+} ∈ Z_{k} if and only if it is of the form

𝔭^{+} = 𝔟′ ⊕ 𝔱^{+}

with 𝔱^{+} maximal isotropic in 𝔱^{ℂ}. Observe that such a 𝔭^{+}is in fact a subalgebra of 𝔤^{ℂ} and so gives rise, by left (or right) translation, to a globally defined Hermitian structure on G. These Hermitian structures were first discovered by Samelson [6] (see also Wang [7]). Once again, we conclude that Z is exhausted by the images of the globally defined Hermitian structures.

One may observe that in both the previous examples there is but a single K-conjugacy class of τ-maximal parabolic subalgebras and thus at most two components of the zero-set of the Nijenhuis tensor. This is a consequence of the fact that both products of spheres and semisimple Lie groups are split-rank symmetric spaces as we now explain.

A symmetric space G/K is said to be split-rank if its rank is the difference between the ranks of G and K. In this case there cannot be any type I_{p} roots since 𝔱_{p} is now maximal abelian in 𝔭. It follows that any τ-stable Borel subalgebra 𝔟 is determined by its 𝔨-part 𝔟_{k} by 𝔟 = 𝔟_{k} +[𝔟_{k}, 𝔱_{p}] and hence that there is a single K-conjugacy class of τ-stable Borel subalgebras. The resulting τ-maximal parabolics built from these Borels by theorem 4.29 of [1] will thus also form a single K-conjugacy class, strengthening proposition 5.4. Since J(𝔱_{p}) has two components, it follows that in the split-rank case there are just one or two components to the zero-set.

Example 3.

Another example of this situation is the symmetric space SU(2n)/Sp(n) which has dimension (2n + 1)(n − 1). This is even for n odd. Let us illustrate the previous sections by determining explicitly the components of its zero-set.

So fix n odd and let V = ℂ^{2n} with its usual Hermitian metric ⟨ , ⟩ and fix a normalized complex volume form ε ∈ Λ^{2n}V^{∗}. A quaternionic structure on V is an antilinear map j: V → V with j^{2} = −1 which is compatible with the metric in the sense that

⟨ju, jv⟩ = ⟨v, u⟩

for all u, v in V. Such a j gives rise to a non-degenerate 2-form ω_{j} ∈ Λ^{2}V^{∗} by

ω_{j}(u, v) = ⟨u, jv⟩

and we further demand that j be compatible with ε in the sense that

ω_{j}^{n} = ε.

Let N be the collection of all such quaternionic structures. Then SU(2n) acts transitively on N by conjugation. A choice of base point j ∈ N allows us to identify ℂ^{2n} with ℍ^{n} and hence its stabilizer with Sp(n). The involution τ corresponding with j is then given by

τ(g) = −jgj

for g ∈ SU(2n). Thus N is a model for SU(2n)/Sp(n).

Fix j and denote ω_{j} by ω. For A ∈ End(V) define the quaternion transpose A^{T} by

ω(Au, v) = ω(u, A^{T} v),

then the symmetric decomposition of the Lie algebra su(2n) is given by

su(2n) = 𝔨 ⊕ 𝔭

with

sp(n) = 𝔨 = {A ∈ su(2n)|A + A^{T} = 0}, 𝔭 = {A ∈ su(2n)|A = A^{T} }.

It is useful to have another model for su(2n) and its complexification. For this we use ω to identify V with V^{∗} by

u(v) = ω(u, v)

so that End(V) ≅ V ⊗ V. Under this identification, it is easy to check that

𝔨^{ℂ} = S^{2}V, 𝔭^{ℂ} = Λ_{0}^{2}V

where Λ_{0}^{2}V is the orthogonal complement of ω ∈ Λ^{2}V^{∗} ≅ Λ^{2}V. Moreover, conjugation with respect to the real form su(2n) becomes

u ⊗ v ↦ jv ⊗ ju

while the involution is given by

u ⊗ v ↦ −(u ⊗ v)^{T} = v ⊗ u.

With these preliminaries, let us fix a maximal torus 𝔱_{k} of 𝔨. This amounts to fixing a j-stable orthogonal decomposition of V into one-dimensional subspaces

V = jL_{n} ⊕ ⋯ ⊕ jL_{1} ⊕ L_{1} ⋯ ⊕ L_{n}.

The fundamental toral subalgebra 𝔱 of su(2n) containing 𝔱_{k} is then the stabilizer in su(2n) of this decomposition.

Any Borel subalgebra 𝔟 of sl(2n, ℂ) containing 𝔱 is the stabilizer of a full flag of subspaces {0} = V_{0} ⊂ V_{1} ⊂ ⋯ ⊂ V_{2n} = V with dim V_{i} = i and each V_{i} a direct sum of some of the L_{i} and jL_{i}. The condition that 𝔟 be τ-stable amounts to the demand that

V_{i}^{0} = V_{2n−i}

where V_{i}^{0} denotes the polar of V_{i} with respect to ω. From this we conclude that, after relabelling the L_{i} if necessary, a τ-stable Borel subalgebra is the stabilizer of a flag given by

V_{i} = ⨁_{k=1}^{i} jL_{n+1−k},  i ≤ n,

V_{n+i} = V_{n−i}^{0} = V_{n} ⊕ ⨁_{k=1}^{i} L_{k},  i ≤ n.

Denoting jL_{k} by L_{−k}, it is now straightforward to check that

𝔟 = ⨁_{i+j≤0} L_{i} ⊗ L_{j}

with nilradical

𝔟′ = ⨁_{i+j<0} L_{i} ⊗ L_{j}.

From 3.5, the τ-maximal parabolic containing 𝔟 has nilradical given by

𝔫 = 𝔟′_{p} + [𝔟′_{p}, 𝔟_{p}].

The following bracketing relations are easy to verify

[L_{i} ∧ L_{−i}, L_{i} ∧ L_{j}] = L_{i} ∨ L_{j},  for i + j < 0, i ≠ j;
[L_{i} ∧ L_{−k}, L_{−i} ∧ L_{−j}] = L_{−k} ∨ L_{−j},  for i ≠ j, k ≥ 0;
[L_{i} ∧ L_{−j}, L_{j} ∧ L_{−k}] = L_{i} ∨ L_{−k},  for 0 < i < j < k;

while all other brackets between summands of 𝔟′_{p} vanish. In particular, S^{2}L_{−k} = [L_{i} ∧ L_{−k}, L_{−i} ∧ L_{−k}] only lies in [𝔟′_{p}, 𝔟_{p}] for 1 ≤ i < k. We therefore conclude that

[𝔟′_{p}, 𝔟_{p}] = ⨁_{\begin{matrix}i+j<0 \\ i≠j \end{matrix}} L_{i} ∨ L_{j} ⊕ ⨁_{i≥2} S^{2}L_{i}.

Thus the 𝔨-part of the Levi factor 𝔩_{k} is given by

𝔩_{k}^{ℂ} = 𝔱_{k}^{ℂ} + S^{2}L_{1} ⊕ S^{2}L_{−1}

so L_{k} = \overbrace{U(1) × ⋯ × U(1)}^{n−1 times} ×Sp(1).

Note that in this case, all τ-stable Borels and hence τ-maximal parabolics are Kconjugate. Thus there are at most two components of the zeroset of the Nijenhuis tensor of J(N), each a copy of the same Z_{𝔮}.

In summary, our analysis shows that any 𝔭^{+} in the zero set arises from an orthogonal decomposition

V = ⨁_{1−n≤i≤n−1} E_{i}

with dim E_{0} = 2, dim E_{i} = 1, |i| ≥ 1 with jE_{i} = E_{−i} (in our previous notation, E_{0} = L_{1} ⊕ L_{−1}, E_{1} = L_{2}, …) and then

𝔭^{+} = ⨁_{i+j<0} E_{i} ∧ E_{j} ⊕ 𝔱^{+}

where 𝔱^{+} is a maximal isotropic subspace of

∑_{i} F_{i} ∧ E_{−i} ∩ {ω}^{⊥}

which is (n − 1)-dimensional.

Example 4. Consideration of example 1 might lead one to enquire as to whether there were Calabi-Eckmann type complex structures on products of odd-dimensional oriented Grassmannians. In fact, this is far from being the case: in this setting, there are, in general, not even any continuous sections of Z as the following theorem shows.

Theorem 6.1. Let M = M_{1} × ⋯ × M_{r} be an even-dimensional product of connected Riemannian symmetric spaces of semisimple type with M_{1} = G_{k}(ℝ^{k+n}) a Grassmannian of oriented k-planes in ℝ^{k+n} with n, k odd and n ≥ k > 1. Then Z has no globally defined continuous sections.

In particular, M admits no Hermitian complex structures.

Proof. A continuous section of Z must lie in some Z_{𝔮} and so gives a reduction of the K-bundle G → G/K = M to some H_{k}. However, τ-maximal parabolic subalgebras commute with the de Rham decomposition of M so that restricting attention to a slice M_{1} ⊂ M, we get a reduction of the SO(k) × SO(n)-bundle SO(n + k) → G_{k}(ℝ^{k+n}) to the centralizer of a maximal torus in SO(k) × SO(n). However, such a reduction would induce a splitting of the tautological k-plane bundle W → G_{k}(ℝ^{k+n}) into a line sub-bundle and its complement:

W = L ⊕ L^{⊥}

and such splittings do not exist for topological reasons. Indeed, such a splitting would give a factorization of Stiefel-Whitney classes

w_{k}(W) = w_{1}(L)w_{k−1}(L^{⊥}),

but w_{1}(L) = 0 since H^{1}(G_{k}(ℝ^{n+k}), ℤ_{2}) vanishes while w_{k}(W) is known to be non-zero, see [4] for example. □

Example 5. Finally, we prove a result of a different nature, relating the topology of G/K to that of the components of Z under the simplifying assumption that K is connected (this involves no loss of generality when G/K is of non-compact type).

We prove

Theorem 6.2. Let G/K be an even-dimensional Riemannian symmetric space of compact or non-compact type with G, K connected and let X be a connected component of Z ⊂ J(G/K). Then

π_{1}(G/K) = π_{1}(X).

Proof. From theorem 5.5, we know that any component of Z arises in the following manner: fix a τ-maximal parabolic subalgebra 𝔮 and let L_{k} be the normalizer of 𝔮 in K. Let 𝔩_{p} be the centralizer of 𝔩_{k} in 𝔭 and take a connected component J_{0}(𝔩_{p}) ⊂ J(𝔩_{p}). Then X = G ×_{K} (K/L_{k} × J_{0}(𝔩_{p})) is a connected component of Z and all components arise this way.

Now L_{k} coincides with the normalizer in K of the parabolic subalgebra 𝔮_{k} of 𝔨^{ℂ} and so is the centralizer of a torus in K whence K/L_{k} is simply connected, as is J_{0}(𝔩_{p}). The homotopy long exact sequence of

K/L_{k} × J_{0}(𝔩_{p}) → X → G/K

now gives

0 → π_{1}(X) → π_{1}(G/K) → π_{0}(K/L_{k} × J_{0}(𝔩_{p})) = 0

whence π_{1}(X) ≅ π_{1}(G/K). □

As a corollary, we see that certain compact quotients of Riemannian symmetric spaces of non-compact type have the same fundamental group as a compact complex manifold. This partially answers a question posed to us by D. Toledo.

Theorem 6.3. Let D be an even-dimensional compact Riemannian locally symmetric space with universal cover M a Riemannian symmetric space of non-compact type. Suppose that, viewed as deck translations, Γ = π_{1}(D) ⊂ I_{0}(M). Then there is a compact complex manifold with fundamental group Γ.

Proof. Let G = I_{0}(M). Then M = G/K with K connected and M simply connected. Let X be a component of Z ⊂ J(M). From theorem 6.2 we know that X is simply connected and, moreover, X is a complex manifold on which G acts holomorphically. Thus Γ\X is the required complex manifold. □

References

[1] F.E. Burstall and J.H. Rawnsley, Twistor theory for Riemannian symmetric spaces, Lecture Notes in Mathematics 1424, Springer-Verlag, Heidelberg, 1990.

[2] E. Calabi and B. Eckmann, A class of compact, complex manifolds which are not algebraic, Ann. Math. 58 (1953), 494-500.

[3] M. Dubois-Violette, Structures complexes au-dessus des variétés, Mathématiques et Physique, Progress in Math. 37, Birkhäuser, Boston, 1983.

[4] John W. Milnor and James D. Stasheff, Characteristic classes, Annals of Mathematics Studies, number 76, Princeton University Press, 1974.

[5] N. O’Brian and J. Rawnsley, Twistor spaces, Annals of Global Analysis and Geometry 3 (1985), 29-58.

[6] H. Samelson, A class of complex-analytic manifolds, Portugal. Math. 12 (1953), 129-132.

[7] H.C. Wang, Closed manifolds with homogeneous complex structure, Amer. J. Math. 76 (1954), 1-32.

School of Mathematics, University of Bath, Bath BA2 7AY, United Kingdom E-mail address: feb@maths.bath.ac.uk

Département de Mathématiques, Université de Metz, Ile du Saulcy, 57045 Metz Cedex, France

E-mail address: gutt@ciril.fr

Mathematics Institute, University of Warwick, Coventry CV4 7AL, United Kingdom E-mail address: jhr@maths.warwick.ac.uk